Home - Search - Browse - Alphabetic Index: 0- 1- 2- 3- 4- 5- 6- 7- 8- 9
A- B- C- D- E- F- G- H- I- J- K- L- M- N- O- P- Q- R- S- T- U- V- W- X- Y- Z
More Details for 2008-01-26
ISS On-Orbit Status 01/26/08

Saturday -- off-duty day for CDR Whitson, FE-1 Malenchenko and FE-2 Tani except for housekeeping and voluntary work.

FE-1 Malenchenko finished Part 2 of his first stress test plus saliva and blood sampling of the ESA/Russian biomed experiment 'IMMUNO' (Neuroendocrine & Immune Responses in Humans During & After Long Term Stay at ISS), today completing remaining urine sample collections. Specimens were then stowed in a special urine containment bag (blood samples were secured yesterday in the MELFI {Minus Eighty Degree Celsius Laboratory Freezer for ISS} in cold packs). (IMMUNO is a 24-hr. test of human immune system changes, with the objective to investigate immune neuro-endocrine reactions in the space environment by studying samples of saliva, blood and urine using collection kits and the biomedical (MBI) protection kit. Also included are entries in a fluid/medications intact log, and a stress-test questionnaire to be filled out by the subject at begin and end of the first day.)

Today CDR Whitson installed the IWIS hardware in preparation for tomorrow's DTF. This will enable a total of 6 Remote Sensor Units (RSU) to collect data for the DTF on Saturday. Remote sensors were installed in the Airlock and Node 2. IWIS is a Station Development Test Objective (SDTO) that records structural measurements to validate math models used in loads analysis. The data will be used to reduce conservatism and might allow relaxation of operational constraints on activities such as crew exercise, vehicle dockings and reboosts that impart structural loads to ISS. The data will also provide more accurate fatigue calculations that could result in an extension of the ISS life.

In support of the end-to-end testing of the ATV PCE (Automated Transfer Vehicle/Proximity Communications Equipment; Russian: MBRL), in the Service Module (SM) Malenchenko installed the ATV control panel (PU), the ATV hand controller, the 'PCE Z0000' box, made the necessary BKS cable connections and also hooked up the BUAP (Antenna Switching Control Box) telemetry connectors to the BITS2-12 onboard telemetry measurement system. The installation work was supported by ground specialist tagup via S-band.

With the Elektron-VM O2 (oxygen) generator currently off, a 5-min cabin air refresh was to be performed by the FE-1 from Progress 27P storage (SrPK) as required.

Starting at ~8:37am EST, TsUP-Moscow conducted five SM thruster tests, each about one second in duration. For the test period, Peggy Whitson had closed the protective Lab science window shutters. The testing concluded at ~8:50am. (Structural dynamics data were taken during the thruster testing by the IWIS (Internal Wireless Instrumentation System), with a total of 6 RSUs (Remote Sensor Units) collecting vibrational data. IWIS is an SDTO (Station Development Test Objective) that records structural measurements to validate math models used in loads analysis. The data will be used to reduce conservatism and might allow relaxation of operational constraints on activities such as crew exercise, vehicle dockings and reboosts that impart structural loads to ISS. The data will also provide more accurate fatigue calculations that could result in an extension of the ISS life.)

In the Airlock, the CDR initiated EMU (Extravehicular Mobility Unit) battery charging in preparation for next week's EVA-14 (1/30).

Malenchenko meanwhile worked on the Russian SRVK-2M condensate water processor, removing its BKO multifiltration/purification unit, which has reached its service life limit, and replacing it with a new unit. The old BKO was stowed for deorbiting in Progress 27. (The BKO, which contains five purification columns to remove dissolved mineral and organic impurities from the condensate, has a service lifetime of at least 450 liters throughput. The purified (deionized) water is used in the Elektron for electrolysis or, after treatment in the BKV water conditioning unit with salts for taste and silver ions for preservation, as potable water in the KPV container.)

For her Voluntary Weekend Science (VolSci) program today, CDR Whitson conducted another session of the LOCAD-PTS (Lab-on-a-Chip Application Development-Portable Test System)/Phase 2 operations, sampling three sites (which showed highest endotoxin levels during the December '07 sampling sessions). Plus two repeats of Negative Control (samples of clean water). (The goal is to compare LOCAD results with the SSK colony growth results. LOCAD uses small, thumb-sized 'microfluidic' cartridges that are read by the experiment reader. The cartridges contain dried extract of horseshoe crab blood cells and colorless dye. In the presence of the bacteria, the dried extract reacts strongly to turn the dye a green color. Therefore, the more green dye, the more microorganisms there are in the original sample. The handheld device tests this new analysis technology by sampling for the presence of gram negative bacteria in the sample in about 15 minutes, showing the results on a display screen. Background: Lab-on-a-Chip technology has an ever-expanding range of applications in the biotech industry. Chips are available (or in development) which can also detect yeast, mold, and gram positive bacteria, identify environmental contaminants, and perform quick health diagnostics in medical clinics. The technology has been used to swab the MERs (Mars Exploration Rovers) for planetary protection. With expanded testing on ISS, began by Sunita Williams in March/April this year, this compact technology has broad potential applications in space exploration--from monitoring environmental conditions to monitoring crew health. The current study should prepare for long-duration exploration by demonstrating a system that enables the crew to perform biochemical analysis in space without having to return samples to Earth.)

The FE-2 connected the ITCS LTL QD (Internal Thermal Control System/Low Temperature Loop/Quick Disconnect) jumper to the CDRA rack (LAB1D6) in support of the activation of the CDRA (Carbon Dioxide Removal Assembly) by MCC-Houston, which requires cooling.

Peggy took air samples for the periodic (currently daily) atmospheric status check for ppO2 (Partial Pressure Oxygen) and ppCO2 (pp Carbon Dioxide), using the hand-held CSA-CP (Compound Specific Analyzer-Combustion Products), CSA-O2 (CSA -Oxygen sensor) and CDMK (CO2 Monitoring Kit). Batteries were to be replaced if necessary. (Purpose of the 15-min activity is to trend with MCA (Major Constituents Analyzer), i.e., to correlate the hand-held readings with MCA measurements.)

Dan Tani performed his daily status check on the BCAT-3 (Binary Colloidal Alloy Test-3) science payload, running by itself in Node-2 since 12/13/07 (briefly interrupted for EVA-13 photo support). (The status check, conducted on the last image taken by the DCS 760 digital still camera which is controlled by EarthKAM software on an A31p laptop, is to verify proper image focus and camera alignment. The SSC (Station Support Computer) is taking photography of the phase separation occurring in the BCAT Sample 3, with the photo flash going off every half hour.)

Dan also completed today's routine maintenance of the SOZh system (Environment Control & Life Support System, ECLSS) in the SM, including ASU toilet facilities systems/replaceables.

The crewmembers performed their regular 2.5-hr physical workout program (about half of which is used for setup & post-exercise personal hygiene) on the TVIS treadmill (CDR, FE-1, FE-2), RED (CDR, FE-2) and VELO bike with bungee cord load trainer (FE-1).

Afterwards, Peggy copied the exercise data file to the MEC (Medical Equipment Computer) laptop for downlink, including the daily wristband HRM (Heart Rate Monitor) data of the workouts on RED, followed by their erasure on the HRM storage medium (done six times a week).

At ~9:00am EST, the crewmembers held their regular weekly planning conference (WPC) with the ground, discussing next week's "Look-Ahead Plan" (prepared jointly by MCC-H and TsUP/Moscow timeline planners) via S-band/audio, reviewing the monthly calendar, upcoming activities, and any concerns about future on-orbit events.

Robotics & SARJ Preview: Continuing the Stbd SARJ (Solar Alpha Rotary Joint) inspection via ground control, today (1/26) the SSRMS (Space Station Remote Manipulator System) was used to inspect the race ring's outer canted surface. Tomorrow, Sunday, the survey will focus on the Datum A surface. During each inspection, the SARJ is rotated using 12 degree directed positioning moves every 5 minutes (requires 8 orbits for a full 360 degree inspection). Meanwhile, ground analysis of the root cause continues. Power analyses for upcoming missions are scheduled to be completed over the next month. SARJ cleaning and lubrication techniques are being evaluated. A grease gun will be added to 1E manifest. ULF-2 manifest will likely have 12 TBAs (Trundle Bearing Assemblies) and 1 DLA (Drive Lock Assembly). Increment 18 and ULF-2 crew will receive training on outboard mode R&R and SARJ cleaning operations.

Weekly Science Update (Expedition Sixteen -- Week 14)

ALTCRISS (Alteino Long Term monitoring of Cosmic Rays on the ISS): Radiation measurements continue to be performed in the PIRS module. Next memory card replacement activity is currently planned on 1/28.

ANITA: Completed.

BCAT-3 (Binary Colloidal Alloy Test 3): 'BCAT-3 Investigator, Peter Lu at Harvard, put together a short movie from the photos taken by astronaut Daniel Tani using EarthKAM to capture the time evolution of sample 3. This captured some great science and showed that the new set-up (improvised by Dan) is working very nicely. Even more excitement occurred when sample 4 began running. Sample 4 is the sample closest to the critical point (on the other side of the critical point) in the phase diagram. It should have a surface tension that is about an order of magnitude less than the previous sample and it was not expected to show structure. BUT, we are seeing something with the new setup, and it is allowing us to see very low contrast images. These new patterns are not yet understood (are they critical fluctuations, bimodal decomposition, something else??). For a short period after sample 4 had been running for about a week and a half some of the images were not good (the camera may have been bumped), but fortunately, the daily camera check and Dan Tani caught this problem. We suspect that he corrected it and we are about to have a look at the new images that have just been downloaded (our camera guru, Jay Owens, will apprise us of the present situation and let everyone know). Because of the new science being seen with sample 4, we have delayed switching to sample 5 until February 4. Three cheers for having astronauts that have the skills and interest to adjust the camera and lighting to capture these difficult to see structures and the interest to check to see that the hardware is working (and right it when it's not).'

CARDIOCOG-2: Completed.

CCISS (Cardiovascular & Cerebrovascular Control on Return from ISS): 'Dan, currently you are scheduled to repeat the last CCISS session on 2/4 and 2/5. The session will need to be repeated due to the Shuttle launch slip moving your last session out of the PI requirements.'

CFE (Capillary Flow Experiment): Reserve.

CSI-2/CGBA (CGBA Science Insert #2/Commercial Generic Bioprocessing Apparatus): In progress.

CGBA-2 (Commercial Generic Bioprocessing Apparatus 2): Complete.

CSLM-2 (Coarsening in Solid-Liquid Mixtures 2): In progress.

EarthKAM (Earth Knowledge Acquired by Middle School Students): Complete.

ELITE-S2 (Elaboratore Immagini Televisive - Space 2): Planned.

EPO (Educational Payload Operations): Reserve.

ETD (Eye Tracking Device): In progress.

IMMUNO (Neuroendocrine & Immune Responses in Humans During & After Long Term Stay at ISS): First session of the experiment for FE-1 Yuri Malenchenko has started on 1/24. Activities continued on 1/25 and 1/26.

InSPACE-2 (Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions 2): In progress.

Integrated Immune: In progress.

KUBIK-FM1/ KUBIK-FM2 Centrifuge/Incubators: Completed.

LOCAD-PTS (Lab-on-a-Chip Application Development-Portable Test System): Complete.

MISSE (Materials ISS Experiment): Ongoing.

MTR-2 (Russian radiation measurements): Passive dosimeters measurements in DC1 'Pirs'.

MULTIGEN-1: MULTIGEN-1 samples will be downloaded on STS-122 (1E).

MSG-SAME (Microgravity Science Glovebox): Complete.

NOA-2 (Nitric Oxide Analyzer): Planned.


PMDIS (Perceptual Motor Deficits in Space): Complete.

SAMS/MAMS (Space & Microgravity Acceleration Measurement Systems): Ongoing.

SLEEP (Sleep-Wake Actigraphy & Light Exposure during Spaceflight): 'Peggy, thanks for downloading and initializing the Actiwatches. We verified the data and will be sending it to the PI. Dan, with your last download activity being completed your only remaining activities are 11 days of Sleep logging and doffing your Actiwatch on the Shuttle with other crewmembers. We are working to add the Sleep logging to the timeline for your convenience.'

SPHERES (Synchronized Position Hold, Engage, Reorient, Experimental Satellite): In progress.

Swab (Characterization of Microorganisms & Allergens in Spacecraft): In progress.

TRAC (Test of Reaction & Adaptation Capabilities): Planned.

ULTRASOUND: 'Dan, the additional Ultrasound troubleshooting this week will be evaluated for a thorough examination of our HRF Rack 1 B1 drawer location. Last Friday's session helped give us a better understanding of the anomaly. We appreciate your patience and interest in assisting in the identification of the anomaly.'

CEO (Crew Earth Observation): Through 1/23 the ground has received a total of 12,027 frames of ISS/CEO imagery for review and cataloging. 'We are pleased to report that your excellent imagery of three of our CEO targets meets or exceeds our current requirements and we can remove them from our standing target list: Luquillo Forest, Puerto Rico; Lake Poopo, Bolivia; and Perth, Australia. Kudos to the crew! Regarding IPY related targets, although we have found none of your atmospheric limb views with polar mesospheric clouds, we have identified your first Aurora imagery acquired over eastern Canada, and, what appears to be three frames of the mega iceberg A53a near South Georgia Island in South Atlantic Ocean. Other fine target area acquisitions under review include those for the South Tibesti Megafans in North Africa and the Patagonian Glaciers. Many of your images of cities at night are absolutely awesome! Your excellent, detailed view of the Port of Suez, Egypt will be published on NASA/GSFC's Earth Observatory website this weekend. You continue to provide us with a wealth of great astronaut photography of the Earth.'

CEO photo targets uplinked for today were Lake Eyre, Australia (weather was predicted to be mostly clear over Lake Eyre. Looking right of track for the lake basin. Context views of the region using the short lens were requested), Patagonian Glaciers, S. America (clear weather conditions continue over the northern Patagonian mountains. Overlapping, nadir mapping frames taken along track of summit glaciers and icefields were requested), and S. Georgia/S. Sandwich, S. Atlantic Ocean (gaps in cloud cover were expected over S. Georgia Island during this overpass. Mapping photography of snow and ice on eastern/southeastern portion of the island was requested).

Home - Search - Browse - Alphabetic Index: 0- 1- 2- 3- 4- 5- 6- 7- 8- 9
A- B- C- D- E- F- G- H- I- J- K- L- M- N- O- P- Q- R- S- T- U- V- W- X- Y- Z
© 1997-2017 Mark Wade - Contact
© / Conditions for Use